per mole undergoes a temperature change of Salve AT = 98 °C - 22°C = 76°C - 10 9 = C, XmX AT = (4.18J/g-K)(250 g)(76 K) = 7.9 I just need help with the practice excersise.

a) Large beds of rocks are used in some solar – heated homes to store heat. Assume that the specific heat of the rocks is 0.82 J/g-K. calculate the quantity of heat absorbed by 50.0 kg of rocks if their temperature increases by 12.0 degrees Celsius. (b) what temperature change would these rocks undergo if they emitted 450 kJ of heat?
Answer #1

Answer -

Given,

Specific Heat of rocks = 0.82 J/g.K

Mass of rocks = 50.0 kg or 50,000 g

°C + 273.15 = K

a) Heat absorbed when Temperature Increase by 12 \degree C = ?

Here the trick is temperature increase in C or K is same. So, temperature increase is 12 K

We know that,

Heat = m * s * \Delta T

Put the values,

Heat absorbed = 50,000 g * 0.82 J/g.K * 12 K

Heat absorbed = 4.92 * 105 J [ANSWER]

b) Temperature change when 450 kJ or 450000 J is emitted = ?

Heat = m * s * \Delta T

Put the values,

450000 J = 50,000 g * 0.82 J/g.K * \Delta T

\DeltaT = (450000 J)/(50,000 g * 0.82 J/g.K)

\DeltaT = 10.9756097561 K or 11 K approx

As the Heat is emitted so Temperature is decreased.

11 K decrease or 11 °C decrease both are same [ANSWER]